### v1

# Contestable AI needs Computational Argumentation

Abstract
AI has become pervasive in recent years, but state-of-the-art approaches predominantly neglect the need for AI systems to be contestable. Instead, contestability is advocated by AI guidelines (e.g. by the OECD) and regulation of automated decision-making (e.g. GDPR). In this position paper we explore how contestability can be achieved computationally in and for AI. We argue that contestable AI requires dynamic (human-machine and/or machine-machine) explainability and decision-making processes, whereby machines can 1. interact with humans and/or other machines to progressively explain their outputs and/or their reasoning as well as assess grounds for contestation provided by these humans and/or other machines, and 2. revise their decision-making processes to redress any issues successfully raised during contestation. Given that much of the current AI landscape is tailored to static AIs, the need to accommodate contestability will require a radical rethinking, that, we argue, computational argumentation is ideally suited to support.
1 Introduction
AI has become pervasive in recent years, with applications ranging from autonomous driving (?) to finance (?) and healthcare (?). Current state-of-the-art AI systems focus on algorithmic solutions, often built from data, generating outputs in the forms of predictions, recommendations and/or decisions and, in some cases, explanations thereof. These existing solutions, though, mostly neglect the need for these AI systems to be contestable. Instead, the need to accommodate contestability is a crucially important problem if AI systems are to be deployed in and benefit society. Indeed, contestability is prominently advocated in some frameworks for AI ethics and in regulations. For example, the Organisation for Economic Co-operation and Development (OECD) states that information should be provided “to enable those adversely affected by an AI system to challenge its outcome based on plain and easy-to-understand information on the factors, and the logic that served as the basis for the prediction, recommendation or decision” (Principle 1.3111https://oecd.ai/en/dashboards/ai-principles/P7.). Moreover, the ACM Global Technology Policy Council for Responsible Algorithmic Systems advocates ‘Contestability and Auditability’ as one of nine principles.222https://www.acm.org/media-center/2022/october/tpc-statement-responsible-algorithmic-systems. Contestability is required by law in some jurisdictions, e.g. GDPR, article 22(3),333https://gdpr-text.com/read/article-22/. states that the data subject’s rights to be safeguarded shall include “at least the right […] to contest the decision”. Finally, contestability and redress are also identified as key principles underpinning the UK AI regulation framework 444https://www.gov.uk/government/publications/ai-regulation-a-pro-innovation-approach/white-paper.
Contestability is seen by some (?) as a means to facilitate accountability, e.g. to prevent the use of Uber’s algorithm for banning drivers because it is not taking into account bias of customers when giving bad reviews. A handful of recent works (overviewed in Section 2) bring contestability to the attention of AI researchers, practitioners and users, while providing insights into its possible interpretations in practice. Moreover, there is evidence, in the form of user studies, that contestability can affect users’ perception of AI fairness (?). However, formal/computational forms of contestability are mostly lacking in the literature, with some exceptions, in the form of 1. indications that, in the case of machine learning, methods generating counterfactual explanations in terms of “actionable changes that an individual can make to flip the prediction of the classifier” may be interpreted as offering (limited) contestability (?) and 2. (?), offering contestability in the very specific setting of causal discovery. Finally, existing works mostly see contestability as a post-hoc process, after predictions/recommendations/decisions have been computed.
In this position paper (specifically in Section 3) we consider what it means for (any) algorithmic decision systems (ADS) (?) to be contestable and contested (by humans or other ADSs). Our take-away message is that formalising/realising computationally contestable AI requires: explanations (by the ADSs which humans/other ADSs may want to contest); grounds for contestation (by the contesters); ability (by the contested) to redress any issues successfully raised during contestation; and ability to interact (by both contested and contester). We then advocate computational argumentation (CA) as being ideally suited to support contestability, providing evidence from the literature on CA (Section 4). We finally conclude (Section 5) with some pointers for future directions in what we believe is a very promising direction of research for the KR community.
2 AI Contestability in the literature
Here, we summarise recent works focusing on advocating or proposing forms of contestability in AI.555 We have identified these works by searching DBLP with ‘Contestable’ and ‘Contestability’, restricting attention to papers on AI from 2014 onwards published in peer-reviewed venues, while also considering additional references therein. We focus on the main contestability dimensions that emerge in this literature, as well as providing pointers to other (less) related research.
Contested entities. Prior work has focused on contesting various aspects of an ADS, including its general design (?; ?; ?), training data (?), training procedure (?), its inputs (?) and outputs (?; ?; ?; ?; ?; ?) as well as the ADS as a whole (?; ?; ?; ?; ?; ?). Note that the possibility of humans being contested by an ADS (e.g., if their assumptions about the subject domain are flawed) is not typically considered.
Contesting entities. All approaches we identified focus on humans as the contesting entities. However, there are considerable differences between the categories and assumed skill sets of people engaged in contestation. Several argue that decision subjects should be empowered to directly contest ADS decisions affecting them, potentially without possessing detailed knowledge of the subject domain (?; ?; ?; ?; ?; ?). Some works suggest that individual decision subjects may not always be able to effectively contest on their own, and propose contestation via third-party representatives (?) or as part of a group in a “class action” (?). Others also consider contestation by professionals, subject matter experts and regulators (?; ?; ?; ?; ?). Note that the possibility of ADSs being the contesting entities is not typically considered.
Contestation methods. Several methods have been proposed for facilitating contestation. In the realm of non-technical solutions, participatory design has been advocated as a way to involve various stakeholders in ADS development, enabling advance mitigation of possible issues and risks associated with the use of the ADS (?; ?; ?). The framework proposed in (?) suggests following a set of development practices over the ADS lifecycle, including incorporating ex-ante safeguards, gathering end-user feedback, implementing quality assurance, mitigating possible risks and allowing for third-party oversight. More technical approaches envision usage of operation logs (?) or model explanations (?; ?; ?; ?; ?). We discuss these and their use of explanations in greater detail next.
The role of explanations. That explanations are needed to support contestability is acknowledged by some (e.g. by (?; ?; ?; ?)), but the problem of using explanations to support contestability in practice has received little attention in explainable AI (XAI) (or AI for that matter) to date. A notable exception is (?), where causal discovery with neural networks guides human feedback for contesting the discovered causal relations, which can be seen as a form of global explanation. Also, some current XAI methods can be seen as offering ground for some limited contestability when they suggest actionable recourse, as is the case for counterfactual explanations in terms of “actionable changes that an individual can make to flip the prediction of the classifier” (?). However, these works disregard that algorithmic decisions may actually be incorrect, e.g. because they are based on incorrect or incomplete data. Moreover, these methods are one-shot, providing no opportunities for follow-up inquiry, and shallow, revealing no information about the steps or logic that led to the explained output, and thus offer little ground for contestability. Some other XAI methods provide information on the AIs’ deliberation and insecurities (?); however, they do not support any form of contestability. Overall, explanations, as understood in state-of-the-art XAI, are seen as inadequate to support contestability (?). Specifically, (?) advocate a generic notion of explanations that capture the rationales behind the development and deployment of the ADS (referred to therein as “process-centric explanations”).
Post-hoc and ex-ante contestability. Some take the view that contestability needs to be supported by regulatory frameworks, involving in particular policy modelling and normative reasoning (?). These approaches see contestability as post-hoc processes, detached from explainability, “to review algorithmic decisions” (?). Ex-ante contestability is envisaged by some, e.g. (?), as a design principle (?), whereby systems are designed to enable users to interactively contest these systems and interaction between humans and systems is needed for “critique and correction” (?), but no technical solutions exist on how to support it.
Contestability as an interaction process. (?) see contestability as an interaction process with humans, at development and deployment times, to allow humans, rather than just data, to train systems. Several other works (e.g. (?; ?)) envisage post-hoc contestability as an interaction process, again between humans and ADSs alone.
Other related work. Works on corrigibility, such as by (?; ?; ?) focus on human overseers providing feedback/instructions to ADSs so that they align with the intentions/values of their user(s) (?; ?). While corrigibility can be viewed as enabling human users to contest the decisions of ADSs, it is narrowly focused on ADSs which optimise for reward when they have incentives to manipulate users (?), ignore instructions (?; ?), and disempower humans/other ADSs (?).
3 An abstract view of AI contestability
In a nutshell, towards supporting contestability, we see ADSs and humans as in Figure 1, assuming that the contested entity is an ADS and the contester is either a human or another ADS.666For simplicity, we assume that the contested ADS is equipped with a single and accompanying and , but in general the same ADS may be equipped with several. Similarly, we assume that the contester is equipped with a single , but in general the same contester may be equipped with several (one per contested model). Further, the same ADS could be both contested and contester.
Let be a model computing outputs in for inputs in . could be. e.g., a machine learning model, an expert system, or a combination of the two. is part of the (contested) ADS, and the outputs of are the ADS’s decisions. For this ADS to be contestable (by a human or another ADS), it needs to be able to process contestations, which may be in reference to one of the following settings:777We ignore here the trivial setting where contestations are directed at specific inputs, as these do not require altering .
(A) output for some specific input , or
(B) how determines from input , or
(C) the full model .
The first setting (A) amounts to contesting the model behaviour without referring to its internal “reasoning” process. The second setting (B) amounts to contesting (part of) the reasoning process of the model for the particular input. The third setting (C) amounts to contesting the model in general, e.g. its “philosophy”, such as the underpinning input and output distributions and/or that it is biased.
Example 1.
As a simple illustration, consider given by a (transparent) set of rules for binary classification amounting to awarding a loan (class loan), or not (class ), with binary features (emp in short) and (breaks in short):888These rules may correspond to the behaviour of a machine learning classifier, where continuous_emp is a “latent feature” of the model. We use them informally here (e.g. they may be implications in classical logic or logic programming rules, with as negation as failure).
For an input (loan applicant) characterised by features emp and breaks, classified as by , in the first setting a contestation may amount to objecting to the classification , whereas in the second setting it may amount to objecting to the instances for of the top and bottom rules (i.e. and ). In the third setting, the middle rule could be the object of contestation, without reference to any specific input.
Explanations for Contestability. In line with some of the literature, we envisage that contestability needs to be supported by explanations for and/or its outputs for specific inputs. We thus assume that the ADS, in addition to , is equipped with an explanation method . Differently from (?), we believe that technical notions of explanations from the literature in XAI may already prove useful. These explanations may be local (i.e. relating to a specific input, output pair) or global (i.e. relating to all input, output pairs). Different forms of contestability may require and/or benefit from different forms of explanations and thus explanation methods, including as follows: 1. a feature-attribution explanation method, e.g. as provided by model-agnostic LIME (?) or SHAP (?) or by model-specific (tailored to neural classifiers) LRP (?) or DeepLIFT (?); 2. an abductive explanation method, e.g. as in (?); 3. a counterfactual explanation method e.g. as first envisaged in (?; ?); 4. a rule-based explanation method e.g as in (?; ?; ?) 5. a mechanistic (local) explanation method, e.g. as provided in DAX (?), SpArX (?), and (?). 6. a mechanistic (global) explanation e.g as in (?); 7. a surrogate model for interpreting , e.g. (?; ?; ?; ?); 8. itself, if already an interpretable model (e.g. a decision tree, or a set of rules).
We stress that satisfying some desirable properties will make the explanation methods more useful to the contestation. Notably, faithfulness requires that the explanations reflect the true reasoning of the explained models (?). This will guide the interactions in contestations towards the correct direction, and allow for more effective redress. Another important property is robustness (?). Non-robust explanation methods could result in drastically different explanations for two users with similar inputs (?), which may jeopardise their explanatory function. In contrast, robust methods (?; ?; ?; ?) improve the consistency and trustworthiness of explanations and may be better suited to support contestability.
Note that, in setting (A), an explanation is not strictly necessary, e.g. a loan applicant could contest not having received a loan no matter the reason. However, if an explanation is present, the contestation can be richer, e.g., if a loan applicant finds out that the reason for the loan refusal is their career-breaks, then they can contest the decision with evidence that they were encouraged to take career-breaks by the employer for training purposes.
Note also that, in settings (B-C), explanations are essential “windows” over the model, without which contestation cannot take place. Specifically, in the second setting, explanations are crucial and, arguably, they need to reveal the reasoning by the model for obtaining the output, e.g. as given by (faithful) mechanistic explanations. Indeed, input-output explanations alone (such as those computed by feature-attribution or counterfactual explanation methods) cannot always provide grounds for contestation. For instance, consider the model in Example 1 and applicant characterised again by features emp and breaks, with the latter due to parental leaves: an explanation including the (instances for of the) first and third rule in the model empowers the loan applicant to contest the decision by objecting to its bias against people who take parental leave; this is more powerful than contesting based on the feature alone in a feature-attribution explanation.
While in the first two settings (A-B) local explanations suffice, in the third setting global explanations are needed. For illustration, a bank manager or regulator, inspecting (interpretable) model in Example 1, may realise of a possible bias underpinning the strict definition of by the first rule, disregarding the possibility of career breaks by applicants.
For further illustration, consider the case of a black-box text classifier which, taken an input text, returns a classification, e.g. the sentiment of the text, or that the text is about a certain topic. Feature-attribution explanations for the classifications (such as those provided by LIME) can be used to pinpoint words in the input text deemed responsible for the classification, but not how the model determines the classification based on those input words – being thus unsuitable for the second and third settings.
Interaction for contestability. The process of contestation needs to be supported by suitable forms of interaction between the contested ADS and the (human or ADS) contester. Interaction may be in the form of conversations in natural language, e.g. using explanations generated by Large Language Models (?), especially if the contester is a human. Alternatively, they may be in structured format, using a formal agent communication language, e.g. in the spirit of FIPA (?).999http://www.fipa.org/. In Figure 1, we indicate with the method used by contested ADS and contester to engage in the interactions necessary for contestability.
The viewpoint of the contester. So far, we have taken the viewpoint of the ADS being contested. The contester may be a human, as in all the existing works on AI contestability, or, alternatively another ADS. In either case, for contestations to be acceptable, they need to be accompanied by some grounds for contestation, as in all earlier illustrations. Thus, we assume that the contester is equipped with a ground generator method . For instance, in setting (A) for Example 1, if a loan applicant finds out that the reason for the loan refusal is their career-breaks and they decide to contest the decision with evidence that they were encouraged to take career-breaks by the employer for training purposes, this evidence forms the grounds for the contestation.
Redress. While explanations empower contestation, its complete realisation needs the ADS to have the ability to redress any issues (successfully) brought about in the contestation. Thus, we assume that the ADS is also equipped with a redress method . For illustration, in the loan application Example 1, may amount to revising the first rule by allowing for exceptions to be made when the breaks are due to training or parental leaves. As an additional illustration for text classifiers, redress may result from the a post-hoc reasoning process with the classifier’s outputs and external knowledge encoded in argument schemes, as in (?), or with the classifier’s outputs and explanations therefor, e.g. as in (?).
Note that we see contestation, and thus redress, as a post-training process, rather than model debugging during training, such as via data augmentation (e.g. as in (?)) or regularisation (e.g. as in (?; ?; ?; ?)). Thus, our view is that contestation is, in general, different from explanation-based model debugging (e.g. as in (?) for tabular data, (?) for text classification, or (?) for image classification). We, however, envisage that, in the case of models trained from data, redress may at times involve fine-tuning or retraining steps in which case the above methods, as well as methods for repairing AI models, e.g. (?; ?; ?), could be drawn upon. In the illustrative case of a black-box text classifier, an example of contestation of the full model (the third setting) is offered by FIND (?): here, LRP (?) is used to associate output neurons in the feature extractor of a text classifier with word clouds. Users can then contest the use of individual neurons by disabling them and fine-tuning the model to no longer rely on them. This allows, specifically, to decrease model bias and reliance on artifacts (?). (?), similarly, aim to expose visual concepts learned by an image classifier using concept activation vectors (?) and fine-tune the model to mitigate the use of specific concepts using gradient regularisation.
4 The Role of Computational Argumentation
We argue that computational argumentation (CA), broadly understood as in (?; ?), is ideally suited to support AI contestability computationally. More strongly, we argue that it is ideally and uniquely suited, as it can cover all aspects of the abstract view in Figure 1 organically. We support this view with reference to several lines of work in the CA literature, as follows.
CA for Explanation. CA has been widely used for XAI (see (?; ?) for recent overviews). It can provide abstractions of several existing, widely used models, e.g. as in (?; ?; ?; ?; ?), and can itself directly serve as the basis of models (?; ?; ?). Natural forms of explanations can be obtained from CA abstractions, e.g. dispute trees (?; ?), defence sets (?) and attribution scores, e.g. gradient-based argument attributions (?), Shapley-based relation attributions (?; ?), amongst several.
CA for Redress. CA provides methods for revising (?; ?; ?) and repairing knowledge bases (?). Much research has been investigated in the context of forgetting (?; ?; ?) and enforcement (?; ?; ?); researchers investigated the effect of expansions (?; ?; ?) and changes in the knowledge base (?; ?; ?; ?). In addition, incomplete argumentation frameworks (?; ?; ?) incorporate uncertainty which enables redress on a conceptual level.
These generic lines of work in CA are useful starting points for supporting redress of models which can be abstracted argumentatively. In addition, reasoning with argumentation frameworks drawn from models is the basis for forms of redress of some (machine learning) models, e.g. as in (?) and (?) for natural language processing, and of some (scheduling) models, as in (?).
CA for Interaction. Interactions between agents, often modelled as dialogues, have been shown to be effectively supported by various forms of CA, e.g. argument schemes (?) or abstract argumentation frameworks (?), in a number of settings. CA’s formal nature allows for principled desiderata of argumentative agent protocols in these settings, as defined in (?). These settings include: computational persuasion (?; ?; ?; ?), framed as selecting the most effective arguments for changing the mind of the other agents; information-seeking and inquiry (?; ?), where agents share information comprising arguments which are private or public; and other areas such as the handling of maliciousness amongst agents (?). Approaches to multi-agent argumentation such as these have been been shown to be useful in various real world applications, e.g. regulatory compliance (?), recommender systems (?; ?; ?; ?) and, more recently, interactive XAI (?; ?; ?). One such approach which looks to have particular promise is that of (?), where argumentative exchanges frame interactive explanation between agents as a conflict resolution problem, while accounting for humans’ cognitive biases. These CA-based approaches to modelling interactions look to have promise for human interaction with Large Language Models, as in (?).
CA and the viewpoint of the contester. CA-based dialogue and interaction methods, e.g. as in the aforementioned (?), already have the potential to accommodate the contester’s viewpoint, leveraging on the use of CA to provide abstractions for underpinning . Indeed, e.g. in (?), contested and contester are interchangeably seen as argumentation frameworks.
5 Conclusion
In this position paper, we have proposed an abstract view of contestable AI and advocated computational argumentation (CA) as ideally positioned to support this view.
Our analysis is restricted, for simplicity and brevity, but could be naturally extended to cover broader scenarios. For example, we have focused on two-parties scenarios where a single ADS is contested and another ADS or a human is doing the contesting; however multi-party scenarios are also possible, e.g. when, in addition to a bank customer, a regulator contests a financial institution. Also, in line with the existing literature on contestability, we have assumed that only ADSs can be contested, but it may be useful to consider the possibility that humans may be contested by ADSs, e.g. in the spirit of (?). In addition to broadening our view to these and other more complex scenarios, future work will be needed to provide evidence of our claim, by building forms of contestable AI that naturally use CA.
Finally, we have advocated CA for contestability, but other KR techniques could provide useful support for some contestability aspects. Specifically, formal verification of AI models could be used to ensure the functional correctness of ADSs, e.g. as in (?; ?). Similarly, rigorous logic-based explainability techniques (?; ?; ?) could be used to ensure the faithfulness and trustworthiness of explanations.
Acknowledgments
This research was partially supported by ERC under the EU’s Horizon 2020 research and innovation programme (grant agreement No. 101020934), by J.P. Morgan and the Royal Academy of Engineering under the Research Chairs and Senior Research Fellowships scheme, by Imperial College through an Imperial College Research Fellowship and by UKRI through the CDT in Safe and Trusted Artificial Intelligence (Grant No. EP/S023356/1) and the CDT in AI for Healthcare (Grant No. EP/S023283/1). Any views or opinions expressed herein are solely those of the authors. We thank Francis Rhys Ward for pointing the authors towards the literature on corrigibility, and Adam Gould, Avinash Kori and Andria Stylianou for helpful discussions.

### v2

@@
 1 Introduction
-AI has become pervasive in recent years, with applications ranging from autonomous driving (?) to finance (?) and healthcare (?).
+AI has become pervasive in recent years, with applications ranging from autonomous driving (Muhammad et al., 2020) to finance (Cao, 2022) and healthcare (Shaheen, 2021).
 Current state-of-the-art AI systems focus on algorithmic solutions, often built from data, generating outputs in the forms of predictions, recommendations and/or decisions and, in some cases, explanations thereof.
@@
 Indeed, contestability is prominently advocated in some frameworks for AI ethics and in regulations.
-For example, the Organisation for Economic Co-operation and Development (OECD) states that information should be provided “to enable those adversely affected by an AI system to challenge its outcome based on plain and easy-to-understand information on the factors, and the logic that served as the basis for the prediction, recommendation or decision” (Principle 1.3111https://oecd.ai/en/dashboards/ai-principles/P7.).
-Moreover, the ACM Global Technology Policy Council for Responsible Algorithmic Systems advocates ‘Contestability and Auditability’ as one of nine principles.222https://www.acm.org/media-center/2022/october/tpc-statement-responsible-algorithmic-systems.
-Contestability is required by law in some jurisdictions, e.g.
-GDPR, article 22(3),333https://gdpr-text.com/read/article-22/.
-states that the data subject’s rights to be safeguarded shall include “at least the right […] to contest the decision”.
-Finally, contestability and redress are also identified as key principles underpinning the UK AI regulation framework 444https://www.gov.uk/government/publications/ai-regulation-a-pro-innovation-approach/white-paper.
-Contestability is seen by some (?) as a means to facilitate accountability, e.g.
+For example, the Organisation for Economic Co-operation and Development (OECD) states that information should be provided “to enable those adversely affected by an AI system to challenge its outcome based on plain and easy-to-understand information on the factors, and the logic that served as the basis for the prediction, recommendation or decision” (Principle 1.3111https://oecd.ai/en/dashboards/ai-principles/P7).
+Moreover, the ACM Global Technology Policy Council for Responsible Algorithmic Systems advocates ‘Contestability and Auditability’ as one of nine principles.222https://www.acm.org/media-center/2022/october/tpc-statement-responsible-algorithmic-systems Contestability is required by law in some jurisdictions, e.g.
+GDPR, article 22(3),333https://gdpr-text.com/read/article-22/ states that the data subject’s rights to be safeguarded shall include “at least the right […] to contest the decision”.
+Finally, contestability and redress are also identified as key principles underpinning the UK AI regulation framework444https://www.gov.uk/government/publications/ai-regulation-a-pro-innovation-approach/white-paper, whereby “Guidance should clarify existing ‘formal’ routes of redress offered by regulators in certain scenarios”.
+This is a major departure point from “classical” requirements for explainable systems, which typically do not require correcting the AI when faulty behaviours are exposed.
+Contestability is seen by some (Hicks, 2022) as a means to facilitate accountability, e.g.
 to prevent the use of Uber’s algorithm for banning drivers because it is not taking into account bias of customers when giving bad reviews.
 A handful of recent works (overviewed in Section 2) bring contestability to the attention of AI researchers, practitioners and users, while providing insights into its possible interpretations in practice.
-Moreover, there is evidence, in the form of user studies, that contestability can affect users’ perception of AI fairness (?).
+Moreover, there is evidence, in the form of user studies, that contestability can affect users’ perception of AI fairness (Yurrita et al., 2023).
 However, formal/computational forms of contestability are mostly lacking in the literature, with some exceptions, in the form of 1.
-indications that, in the case of machine learning, methods generating counterfactual explanations in terms of “actionable changes that an individual can make to flip the prediction of the classifier” may be interpreted as offering (limited) contestability (?) and 2.
-(?), offering contestability in the very specific setting of causal discovery.
+indications that, in the case of machine learning, methods generating counterfactual explanations in terms of “actionable changes that an individual can make to flip the prediction of the classifier” may be interpreted as offering (limited) contestability (Venkatasubramanian and Alfano, 2020) and 2.
+(Russo and Toni, 2023), offering contestability in the very specific setting of causal discovery.
 Finally, existing works mostly see contestability as a post-hoc process, after predictions/recommendations/decisions have been computed.
-In this position paper (specifically in Section 3) we consider what it means for (any) algorithmic decision systems (ADS) (?) to be contestable and contested (by humans or other ADSs).
+In this position paper (specifically in Section 3) we consider what it means for (any) algorithmic decision systems (ADS) (Henin and Métayer, 2022) to be contestable and contested (by humans or other ADSs).
 Our take-away message is that formalising/realising computationally contestable AI requires: explanations (by the ADSs which humans/other ADSs may want to contest); grounds for contestation (by the contesters); ability (by the contested) to redress any issues successfully raised during contestation; and ability to interact (by both contested and contester).
@@
 Contested entities.
-Prior work has focused on contesting various aspects of an ADS, including its general design (?; ?; ?), training data (?), training procedure (?), its inputs (?) and outputs (?; ?; ?; ?; ?; ?) as well as the ADS as a whole (?; ?; ?; ?; ?; ?).
+Prior work has focused on contesting various aspects of an ADS, including its general design (Almada, 2019; Alfrink et al., 2022; Yurrita, Balayn, and Gadiraju, 2023), training data (Kluttz, Kohli, and Mulligan, 2022), training procedure (Almada, 2019), its inputs (Alfrink et al., 2022) and outputs (Hirsch et al., 2017; Ploug and Holm, 2020; Tubella et al., 2020; Lyons, Velloso, and Miller, 2021; Henin and Métayer, 2022; Hicks, 2022) as well as the ADS as a whole (Lyons, Velloso, and Miller, 2021; Vaccaro et al., 2021; Henin and Métayer, 2022; Hicks, 2022; Alfrink et al., 2023; Russo and Toni, 2023).
 Note that the possibility of humans being contested by an ADS (e.g., if their assumptions about the subject domain are flawed) is not typically considered.
@@
 However, there are considerable differences between the categories and assumed skill sets of people engaged in contestation.
-Several argue that decision subjects should be empowered to directly contest ADS decisions affecting them, potentially without possessing detailed knowledge of the subject domain (?; ?; ?; ?; ?; ?).
-Some works suggest that individual decision subjects may not always be able to effectively contest on their own, and propose contestation via third-party representatives (?) or as part of a group in a “class action” (?).
-Others also consider contestation by professionals, subject matter experts and regulators (?; ?; ?; ?; ?).
+Several argue that decision subjects should be empowered to directly contest ADS decisions affecting them, potentially without possessing detailed knowledge of the subject domain (Almada, 2019; Ploug and Holm, 2020; Lyons, Velloso, and Miller, 2021; Vaccaro et al., 2021; Henin and Métayer, 2022; Hicks, 2022).
+Some works suggest that individual decision subjects may not always be able to effectively contest on their own, and propose contestation via third-party representatives (Alfrink et al., 2022) or as part of a group in a “class action” (Lyons, Velloso, and Miller, 2021).
+Others also consider contestation by professionals, subject matter experts and regulators (Tubella et al., 2020; Henin and Métayer, 2022; Kluttz, Kohli, and Mulligan, 2022; Alfrink et al., 2023; Russo and Toni, 2023).
 Note that the possibility of ADSs being the contesting entities is not typically considered.
@@
 Several methods have been proposed for facilitating contestation.
-In the realm of non-technical solutions, participatory design has been advocated as a way to involve various stakeholders in ADS development, enabling advance mitigation of possible issues and risks associated with the use of the ADS (?; ?; ?).
-The framework proposed in (?) suggests following a set of development practices over the ADS lifecycle, including incorporating ex-ante safeguards, gathering end-user feedback, implementing quality assurance, mitigating possible risks and allowing for third-party oversight.
-More technical approaches envision usage of operation logs (?) or model explanations (?; ?; ?; ?; ?).
+In the realm of non-technical solutions, participatory design has been advocated as a way to involve various stakeholders in ADS development, enabling advance mitigation of possible issues and risks associated with the use of the ADS (Almada, 2019; Vaccaro et al., 2021; Alfrink et al., 2023).
+The framework proposed in (Alfrink et al., 2022) suggests following a set of development practices over the ADS lifecycle, including incorporating ex-ante safeguards, gathering end-user feedback, implementing quality assurance, mitigating possible risks and allowing for third-party oversight.
+More technical approaches envision usage of operation logs (Hicks, 2022) or model explanations (Almada, 2019; Ploug and Holm, 2020; Kluttz, Kohli, and Mulligan, 2022; Russo and Toni, 2023; Yurrita, Balayn, and Gadiraju, 2023).
 We discuss these and their use of explanations in greater detail next.
@@
 That explanations are needed to support contestability is acknowledged by some (e.g.
-by (?; ?; ?; ?)), but the problem of using explanations to support contestability in practice has received little attention in explainable AI (XAI) (or AI for that matter) to date.
-A notable exception is (?), where causal discovery with neural networks guides human feedback for contesting the discovered causal relations, which can be seen as a form of global explanation.
-Also, some current XAI methods can be seen as offering ground for some limited contestability when they suggest actionable recourse, as is the case for counterfactual explanations in terms of “actionable changes that an individual can make to flip the prediction of the classifier” (?).
+by Almada (2019); Lyons, Velloso, and Miller (2021); Alfrink et al.
+(2022); Wachter, Mittelstadt, and Russell (2017)), but the problem of using explanations to support contestability in practice has received little attention in explainable AI (XAI) (or AI for that matter) to date.
+A notable exception is (Russo and Toni, 2023), where causal discovery with neural networks guides human feedback for contesting the discovered causal relations, which can be seen as a form of global explanation.
+Also, some current XAI methods can be seen as offering ground for some limited contestability when they suggest actionable recourse, as is the case for counterfactual explanations in terms of “actionable changes that an individual can make to flip the prediction of the classifier” (Venkatasubramanian and Alfano, 2020).
 However, these works disregard that algorithmic decisions may actually be incorrect, e.g.
@@
 Moreover, these methods are one-shot, providing no opportunities for follow-up inquiry, and shallow, revealing no information about the steps or logic that led to the explained output, and thus offer little ground for contestability.
-Some other XAI methods provide information on the AIs’ deliberation and insecurities (?); however, they do not support any form of contestability.
-Overall, explanations, as understood in state-of-the-art XAI, are seen as inadequate to support contestability (?).
-Specifically, (?) advocate a generic notion of explanations that capture the rationales behind the development and deployment of the ADS (referred to therein as “process-centric explanations”).
+Some other XAI methods provide information on the AIs’ deliberation and insecurities (Wang and Vasconcelos, 2019); however, they do not support any form of contestability.
+Overall, explanations, as understood in state-of-the-art XAI, are seen as inadequate to support contestability (Henin and Métayer, 2022).
+Specifically, Yurrita, Balayn, and Gadiraju (2023) advocate a generic notion of explanations that capture the rationales behind the development and deployment of the ADS (referred to therein as “process-centric explanations”).
 Post-hoc and ex-ante contestability.
-Some take the view that contestability needs to be supported by regulatory frameworks, involving in particular policy modelling and normative reasoning (?).
-These approaches see contestability as post-hoc processes, detached from explainability, “to review algorithmic decisions” (?).
-Ex-ante contestability is envisaged by some, e.g.
-(?), as a design principle (?), whereby systems are designed to enable users to interactively contest these systems and interaction between humans and systems is needed for “critique and correction” (?), but no technical solutions exist on how to support it.
+Some take the view that contestability needs to be supported by regulatory frameworks, involving in particular policy modelling and normative reasoning (Tubella et al., 2020).
+These approaches see contestability as post-hoc processes, detached from explainability, “to review algorithmic decisions” (Lyons, Velloso, and Miller, 2021).
+Ex-ante contestability is envisaged by some as a design principle (Hirsch et al., 2017; Alfrink et al., 2022), whereby systems are designed to enable users to interactively contest these systems and interaction between humans and systems is needed for “critique and correction” (Lyons, Velloso, and Miller, 2021), but no technical solutions exist on how to support it.
 Contestability as an interaction process.
-(?) see contestability as an interaction process with humans, at development and deployment times, to allow humans, rather than just data, to train systems.
+Kluttz, Kohli, and Mulligan (2022) see contestability as an interaction process with humans, at development and deployment times, to allow humans, rather than just data, to train systems.
 Several other works (e.g.
-(?; ?)) envisage post-hoc contestability as an interaction process, again between humans and ADSs alone.
+(Hirsch et al., 2017; Lyons, Velloso, and Miller, 2021)) envisage post-hoc contestability as an interaction process, again between humans and ADSs alone.
 Other related work.
-Works on corrigibility, such as by (?; ?; ?) focus on human overseers providing feedback/instructions to ADSs so that they align with the intentions/values of their user(s) (?; ?).
-While corrigibility can be viewed as enabling human users to contest the decisions of ADSs, it is narrowly focused on ADSs which optimise for reward when they have incentives to manipulate users (?), ignore instructions (?; ?), and disempower humans/other ADSs (?).
+Works on corrigibility, such as (Soares et al., 2015; Carey and Everitt, 2023; Russell, Dewey, and Tegmark, 2015) focus on human overseers providing feedback/instructions to ADSs so that they align with the intentions/values of their user(s) (Gabriel, 2020; Hadfield-Menell et al., 2016).
+While corrigibility can be viewed as enabling human users to contest the decisions of ADSs, it is narrowly focused on ADSs which optimise for reward when they have incentives to manipulate users (Ward, Toni, and Belardinelli, 2022), ignore instructions (Hadfield-Menell et al., 2016), and disempower humans/other ADSs (Turner et al., 2021).
 3 An abstract view of AI contestability
@@
 We thus assume that the ADS, in addition to , is equipped with an explanation method .
-Differently from (?), we believe that technical notions of explanations from the literature in XAI may already prove useful.
+Differently from Yurrita, Balayn, and Gadiraju (2023), we believe that technical notions of explanations from the literature in XAI may already prove useful.
 These explanations may be local (i.e.
@@
 a feature-attribution explanation method, e.g.
-as provided by model-agnostic LIME (?) or SHAP (?) or by model-specific (tailored to neural classifiers) LRP (?) or DeepLIFT (?); 2.
+as provided by model-agnostic LIME (Ribeiro, Singh, and Guestrin, 2016) or SHAP (Lundberg and Lee, 2017) or by model-specific (tailored to neural classifiers) LRP (Montavon et al., 2019) or DeepLIFT (Shrikumar, Greenside, and Kundaje, 2017); 2.
 an abductive explanation method, e.g.
-as in (?); 3.
+as in (Ignatiev, Narodytska, and Marques-Silva, 2019); 3.
 a counterfactual explanation method e.g.
-as first envisaged in (?; ?); 4.
-a rule-based explanation method e.g as in (?; ?; ?) 5.
+as first envisaged in (Wachter, Mittelstadt, and Russell, 2017; Tolomei et al., 2017); 4.
+a rule-based explanation method e.g as in (Ribeiro, Singh, and Guestrin, 2018; Guidotti et al., 2019; Ferreira et al., 2022) 5.
 a mechanistic (local) explanation method, e.g.
-as provided in DAX (?), SpArX (?), and (?).
+as provided in DAX (Albini et al., 2020), SpArX (Ayoobi, Potyka, and Toni, 2023), and (Wang et al., 2023).
 6.
-a mechanistic (global) explanation e.g as in (?); 7.
+a mechanistic (global) explanation e.g as in (Ayoobi, Potyka, and Toni, 2023); 7.
 a surrogate model for interpreting , e.g.
-(?; ?; ?; ?); 8.
+(Kenny and Keane, 2019; Tan and Kotthaus, 2022; Potyka, Yin, and Toni, 2023); 8.
 itself, if already an interpretable model (e.g.
@@
 We stress that satisfying some desirable properties will make the explanation methods more useful to the contestation.
-Notably, faithfulness requires that the explanations reflect the true reasoning of the explained models (?).
+Notably, faithfulness requires that the explanations reflect the true reasoning of the explained models (Lakkaraju et al., 2019).
 This will guide the interactions in contestations towards the correct direction, and allow for more effective redress.
-Another important property is robustness (?).
-Non-robust explanation methods could result in drastically different explanations for two users with similar inputs (?), which may jeopardise their explanatory function.
-In contrast, robust methods (?; ?; ?; ?) improve the consistency and trustworthiness of explanations and may be better suited to support contestability.
+Another important property is robustness (Jiang et al., 2024b).
+Non-robust explanation methods could result in drastically different explanations for two users with similar inputs (Leofante and Potyka, 2024), which may jeopardise their explanatory function.
+In contrast, robust methods (Jiang et al., 2023; Leofante and Lomuscio, 2023; Leofante, Botoeva, and Rajani, 2023; Jiang et al., 2024a) improve the consistency and trustworthiness of explanations and may be better suited to support contestability.
 Note that, in setting (A), an explanation is not strictly necessary, e.g.
@@
 Interaction may be in the form of conversations in natural language, e.g.
-using explanations generated by Large Language Models (?), especially if the contester is a human.
+using explanations generated by Large Language Models (Bills et al., 2023), especially if the contester is a human.
 Alternatively, they may be in structured format, using a formal agent communication language, e.g.
-in the spirit of FIPA (?).999http://www.fipa.org/.
-In Figure 1, we indicate with the method used by contested ADS and contester to engage in the interactions necessary for contestability.
+in the spirit of FIPA (Poslad, 2007).999http://www.fipa.org/ In Figure 1, we indicate with the method used by contested ADS and contester to engage in the interactions necessary for contestability.
 The viewpoint of the contester.
@@
 For illustration, in the loan application Example 1, may amount to revising the first rule by allowing for exceptions to be made when the breaks are due to training or parental leaves.
-As an additional illustration for text classifiers, redress may result from the a post-hoc reasoning process with the classifier’s outputs and external knowledge encoded in argument schemes, as in (?), or with the classifier’s outputs and explanations therefor, e.g.
-as in (?).
+As an additional illustration for text classifiers, redress may result from the a post-hoc reasoning process with the classifier’s outputs and external knowledge encoded in argument schemes, as in (Carstens and Toni, 2017), or with the classifier’s outputs and explanations therefor, e.g.
+as in (Freedman et al., 2024).
 Note that we see contestation, and thus redress, as a post-training process, rather than model debugging during training, such as via data augmentation (e.g.
-as in (?)) or regularisation (e.g.
-as in (?; ?; ?; ?)).
+as in (Teso and Kersting, 2019)) or regularisation (e.g.
+as in (Ross, Hughes, and Doshi-Velez, 2017; Rieger et al., 2020; Shao et al., 2021; Zhang, Williams, and Toni, 2024)).
 Thus, our view is that contestation is, in general, different from explanation-based model debugging (e.g.
-as in (?) for tabular data, (?) for text classification, or (?) for image classification).
+as in (Ghai et al., 2021) for tabular data, (Lertvittayakumjorn and Toni, 2021) for text classification, or (Popordanoska, Kumar, and Teso, 2020) for image classification).
 We, however, envisage that, in the case of models trained from data, redress may at times involve fine-tuning or retraining steps in which case the above methods, as well as methods for repairing AI models, e.g.
-(?; ?; ?), could be drawn upon.
-In the illustrative case of a black-box text classifier, an example of contestation of the full model (the third setting) is offered by FIND (?): here, LRP (?) is used to associate output neurons in the feature extractor of a text classifier with word clouds.
+(Henriksen, Leofante, and Lomuscio, 2022; Almog and Kalech, 2023), could be drawn upon.
+In the illustrative case of a black-box text classifier, an example of contestation of the full model (the third setting) is offered by FIND (Lertvittayakumjorn, Specia, and Toni, 2020): here, LRP (Montavon et al., 2019) is used to associate output neurons in the feature extractor of a text classifier with word clouds.
 Users can then contest the use of individual neurons by disabling them and fine-tuning the model to no longer rely on them.
-This allows, specifically, to decrease model bias and reliance on artifacts (?).
-(?), similarly, aim to expose visual concepts learned by an image classifier using concept activation vectors (?) and fine-tune the model to mitigate the use of specific concepts using gradient regularisation.
+This allows, specifically, to decrease model bias and reliance on artifacts (Lertvittayakumjorn, Specia, and Toni, 2020).
+Dreyer et al.
+(2024), similarly, aim to expose visual concepts learned by an image classifier using concept activation vectors (Kim et al., 2018) and fine-tune the model to mitigate the use of specific concepts using gradient regularisation.
 4 The Role of Computational Argumentation
-We argue that computational argumentation (CA), broadly understood as in (?; ?), is ideally suited to support AI contestability computationally.
-More strongly, we argue that it is ideally and uniquely suited, as it can cover all aspects of the abstract view in Figure 1 organically.
+In this position paper we argue that computational argumentation (CA) is ideally suited to support AI contestability computationally.
+CA, broadly understood as in (Atkinson et al., 2017; Baroni et al., 2018), is a branch of Knowledge Representation & Reasoning which represents information in terms of arguments and dialectical relations (of attack and, possibly, support) between them.
+CA is equipped with semantics to reach some form of consensus regarding conclusions to be drawn.
+As such, CA is ideally and uniquely suited to cover all aspects of the abstract view in Figure 1 organically.
 We support this view with reference to several lines of work in the CA literature, as follows.
 CA for explanation.
-CA has been widely used for XAI (see (?; ?) for recent overviews).
+CA has been widely used for XAI (see (Vassiliades, Bassiliades, and Patkos, 2021; Cyras et al., 2021) for recent overviews).
 It can provide abstractions of several existing, widely used models, e.g.
-as in (?; ?; ?; ?; ?), and can itself directly serve as the basis of models (?; ?; ?).
+as in (Potyka, Yin, and Toni, 2023; Ayoobi, Potyka, and Toni, 2023; Prakken and others, 2020; Fan, 2018; Cyras et al., 2019b), and can itself directly serve as the basis of models (Rago, Cocarascu, and Toni, 2018; Rago et al., 2020; Cocarascu et al., 2020).
 Natural forms of explanations can be obtained from CA abstractions, e.g.
-dispute trees (?; ?), defence sets (?) and attribution scores, e.g.
-gradient-based argument attributions (?), Shapley-based relation attributions (?; ?), amongst several.
+dispute trees (Fan and Toni, 2015b; Čyras et al., 2019a), defence sets (Arioua, Tamani, and Croitoru, 2015) and attribution scores, e.g.
+gradient-based argument attributions (Yin, Potyka, and Toni, 2023), Shapley-based relation attributions (Amgoud, Ben-Naim, and Vesic, 2017; Yin, Potyka, and Toni, 2024), amongst several.
 CA for redress.
-CA provides methods for revising (?; ?; ?) and repairing knowledge bases (?).
-Much research has been investigated in the context of forgetting (?; ?; ?) and enforcement (?; ?; ?); researchers investigated the effect of expansions (?; ?; ?) and changes in the knowledge base (?; ?; ?; ?).
-In addition, incomplete argumentation frameworks (?; ?; ?) incorporate uncertainty which enables redress on a conceptual level.
+CA provides methods for revising (Snaith and Reed, 2017; Baumann and Brewka, 2015; Falappa, Kern-Isberner, and Simari, 2009) and repairing knowledge bases (Ulbricht and Baumann, 2019).
+Much research has been investigated in the context of forgetting (Berthold, Rapberger, and Ulbricht, 2023; Baumann, Gabbay, and Rodrigues, 2020) and enforcement (Rapberger and Ulbricht, 2023; Baumann et al., 2021); researchers investigated the effect of expansions (Prakken, 2023; Oikarinen and Woltran, 2011; Cayrol, de Saint-Cyr, and Lagasquie-Schiex, 2010) and changes in the knowledge base (Doutre and Mailly, 2018; Booth et al., 2013; Moguillansky et al., 2013; Niskanen, 2020).
+In addition, incomplete argumentation frameworks (Baumeister et al., 2021; Alfano et al., 2023) incorporate uncertainty which enables redress on a conceptual level.
 These generic lines of work in CA are useful starting points for supporting redress of models which can be abstracted argumentatively.
 In addition, reasoning with argumentation frameworks drawn from models is the basis for forms of redress of some (machine learning) models, e.g.
-as in (?) and (?) for natural language processing, and of some (scheduling) models, as in (?).
+as in (Carstens and Toni, 2017) and (Freedman et al., 2024) for natural language processing, and of some (scheduling) models, as in (Cyras et al., 2019b).
 CA for interaction.
 Interactions between agents, often modelled as dialogues, have been shown to be effectively supported by various forms of CA, e.g.
-argument schemes (?) or abstract argumentation frameworks (?), in a number of settings.
-CA’s formal nature allows for principled desiderata of argumentative agent protocols in these settings, as defined in (?).
-These settings include: computational persuasion (?; ?; ?; ?), framed as selecting the most effective arguments for changing the mind of the other agents; information-seeking and inquiry (?; ?), where agents share information comprising arguments which are private or public; and other areas such as the handling of maliciousness amongst agents (?).
-Approaches to multi-agent argumentation such as these have been been shown to be useful in various real world applications, e.g.
-regulatory compliance (?), recommender systems (?; ?; ?; ?) and, more recently, interactive XAI (?; ?; ?).
-One such approach which looks to have particular promise is that of (?), where argumentative exchanges frame interactive explanation between agents as a conflict resolution problem, while accounting for humans’ cognitive biases.
-These CA-based approaches to modelling interactions look to have promise for human interaction with Large Language Models, as in (?).
+argument schemes (Panisson, McBurney, and Bordini, 2021) or abstract argumentation frameworks (de Tarlé, Bonzon, and Maudet, 2022), in a number of settings.
+CA’s formal nature allows for principled desiderata of argumentative agent protocols in these settings, as defined in (McBurney, Parsons, and Wooldridge, 2002).
+These settings include: computational persuasion (Fan and Toni, 2012; Hunter, 2018; Calegari, Riveret, and Sartor, 2021; Donadello et al., 2022), framed as selecting the most effective arguments for changing the mind of the other agents; information-seeking and inquiry (Black and Hunter, 2007; Fan and Toni, 2015a), where agents share information comprising arguments which are private or public; and other areas such as the handling of maliciousness amongst agents (Kontarinis and Toni, 2015).
+Approaches to multi-agent argumentation such as these have been been shown to be useful in various real-world applications, e.g.
+regulatory compliance (Raymond et al., 2022), recommender systems (Briguez et al., 2014; Teze, Godo, and Simari, 2018; Rago et al., 2020, 2021) and, more recently, interactive XAI (Madumal et al., 2019; Calegari et al., 2022; Paulino-Passos and Toni, 2022).
+One such approach which looks to have particular promise is that of Rago, Li, and Toni (2023), where argumentative exchanges frame interactive explanation between agents as a conflict resolution problem, while accounting for humans’ cognitive biases.
+These CA-based approaches to modelling interactions look to have promise for human interaction with Large Language Models, as in (Freedman et al., 2024).
 CA and the viewpoint of the contester.
 CA-based dialogue and interaction methods, e.g.
-as in the aforementioned (?), already have the potential to accommodate the contester’s viewpoint, leveraging on the use of CA to provide abstractions for underpinning .
+as in the aforementioned (Rago, Li, and Toni, 2023), already have the potential to accommodate the contester’s viewpoint, leveraging on the use of CA to provide abstractions for underpinning .
 Indeed, e.g.
-in (?), contested and contester are interchangeably seen as argumentation frameworks.
-5 Conclusion
+in (Rago, Li, and Toni, 2023), contested and contester are interchangeably seen as argumentation frameworks.
+5 Discussion and Future Work
 In this position paper, we have proposed an abstract view of contestable AI and advocated computational argumentation (CA) as ideally positioned to support this view.
 Our analysis is restricted, for simplicity and brevity, but could be naturally extended to cover broader scenarios.
-For example, we have focused on two-parties scenarios where a single ADS is contested and another ADS or a human is doing the contesting; however multi-party scenarios are also possible, e.g.
+For example, we have focused on two-party scenarios where a single ADS is contested and another ADS or a human is doing the contesting; however multi-party scenarios are also possible, e.g.
 when, in addition to a bank customer, a regulator contests a financial institution.
 Also, in line with the existing literature on contestability, we have assumed that only ADSs can be contested, but it may be useful to consider the possibility that humans may be contested by ADSs, e.g.
-in the spirit of (?).
+in the spirit of (Miller, 2023).
 In addition to broadening our view to these and other more complex scenarios, future work will be needed to provide evidence of our claim, by building forms of contestable AI that naturally use CA.
+This involves connecting what we see are the main requirements of contestability (explanations, grounds for contestation, redress, and interaction) in concrete applications and scenarios.
+CA has already been used to solve several individual challenges, e.g.
+see (Borg and Bex, 2020; Lawrence, Visser, and Reed, 2023; Cyras et al., 2019b).
+However, considerable engineering work is needed to combine solutions towards fully-fledged end-to-end contestable systems.
+For instance, in order to support contestability with modern AI systems, we envisage that considerable effort will be required to build large-scale argumentation solvers.
 Finally, we have advocated CA for contestability, but other KR techniques could provide useful support for some contestability aspects.
 Specifically, formal verification of AI models could be used to ensure the functional correctness of ADSs, e.g.
-as in (?; ?).
-Similarly, rigorous logic-based explainability techniques (?; ?; ?) could be used to ensure the faithfulness and trustworthiness of explanations.
+as in (Kouvaros et al., 2023).
+Similarly, rigorous logic-based explainability techniques (Narodytska et al., 2019; Darwiche, 2020) could be used to ensure the faithfulness and trustworthiness of explanations.
 Acknowledgments